A gas X is collected over water at 17 degree C and 750 mm pressure. If the volume of the gas collected is 50 cc, calculate the volume of the dry gas at s.t.p. (at 17degree C the vapour pressure is 14mm)

Grade:9

1 Answers

Ravleen Kaur
6 years ago
Hello student,

Given,
Pressure initially, P1= 750 mm
Pressure finally, P2 = 14 mm
Temperature , T1 = T2= 170C =17+273 = 290 K
Volume Initially V1= 50 cc
Formula Used:
P1V1/T1 =P2V2/T2
Volume finally, V2=P1V1T2/P2T1= (750 X 50)/14 = 2678.5714 cc

Regards
IHE